OpenStudy (anonymous):

how large should n be so that the midpoint rule approximation of definite integral 1 / x on [1,2] is accurate to within 10^-3 . hint | Em| < = K (b-a)^3 / (24n^2) where |f '' (x) | < = k on [ a, b]

OpenStudy (anonymous):

the second derive is 2/x^3, which is decreasing on [ 1, 2 ]

OpenStudy (anonymous):

Man, instead of counting rectangles, ask your teacher why you should count rectangles when you won't need to again after a week when you are introduced to integrals

OpenStudy (anonymous):

so | f '' (x) | < 2/1^3

OpenStudy (anonymous):

you do need to when you cant find the antiderivative

OpenStudy (anonymous):

like integral e^(-x^2) or integral sin (x^2)

OpenStudy (anonymous):

That's when you reach for your handy TI-83, since it is an approximation anyway...

OpenStudy (anonymous):

when you want the definite integral

OpenStudy (anonymous):

suppose you work in industry and they need more precision than a TI 83

OpenStudy (anonymous):

you want to bound your error

OpenStudy (anonymous):

midpoint formula is not more precise than TI-83

OpenStudy (anonymous):

yes it is , if n is large enough

OpenStudy (anonymous):

Why not Simpson's rule, then?

OpenStudy (anonymous):

i have no clue what the precision of TI 83 is, but i know the error of midpoint approximation

OpenStudy (anonymous):

they all work, im just using midpoint at the moment

OpenStudy (anonymous):

you asked why not just use TI 83, because TI 83 cannot be made as precise as you want. it has a limit, which i dont even know what it is , what decimal

OpenStudy (anonymous):

how precise to what decimal

OpenStudy (anonymous):

9, I think

OpenStudy (anonymous):

see what i mean

OpenStudy (anonymous):

you have no idea. thats why we have those exact bound formulas , no guess work here

OpenStudy (anonymous):

maybe 10 , maybe 8 , who knows?

OpenStudy (anonymous):

Ti 83 is fine for quick work, not for sensitive industrial stuff

OpenStudy (anonymous):

Okay, so how do you do this?

OpenStudy (anonymous):

i dont know, in math we like to talk about making the error as small as we please, thats important because you want or expect the error to go to zero . if the error does not go to zero then there is no definitive answer (divergent)

OpenStudy (anonymous):

2/x^-3 on [1,2] = 2 So K<2

OpenStudy (anonymous):

and K (b-a)^3 / (24n^2) = 2(1)/(24n^2) = 10^-3 --> 10^3/12 = n^2 n = sqrt(250/3) = 9.128... so 10

OpenStudy (anonymous):

no, that would make the error more precise than it actually is